OpenStudy (ally_b):

Can anyone help me answer this question? Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (phi):

do you know how to find the "length" or magnitude of the vector u ?

OpenStudy (ally_b):

No

OpenStudy (phi):

do you know how to do a "dot product" of u * u ?

OpenStudy (ally_b):

yes

OpenStudy (phi):

the magnitude of u is sqrt(u dot u) first, what is u dot u = ?

OpenStudy (ally_b):

I have no idea... I am confused

OpenStudy (phi):

the dot product of two vectors <a,b> and <c,d> is a*c + b*d

OpenStudy (phi):

if it is a vector dotted with itself, for example <a , b> dot <a, b> you get a*a + b*b

OpenStudy (phi):

can you do <-4, -3> dot <-4, -3 > ?

OpenStudy (ally_b):

So 16+9=25?

OpenStudy (phi):

yes and the length of vector u is sqrt (u dot u) so sqr(25)

OpenStudy (phi):

sqr(25) simplifies

OpenStudy (ally_b):

so 5

OpenStudy (phi):

to make vector u "unit length", divide each of its "components" by the length of u in other words divide -4 and -3 by 5

OpenStudy (ally_b):

so it would be -4-3/5 right?

OpenStudy (ally_b):

-4/5, -3/5

OpenStudy (phi):

yes, and if you find the length of that vector you will find it is 1 sqrt( <-4/5, -3/5> dot < -4/5, -3/5> ) = sqrt( 16/25 + 9/25) = sqr( 25/25) = sqr(1) = 1

OpenStudy (ally_b):

Thank you so much!!